# Establishing Best Practices for Building Rigorous Agentic Benchmarks

### v4

@@ 1 Introduction
Although issues in evaluation rigor can significantly skew evaluation results, they are still frequently overlooked in the current development, deployment, and analysis of agentic benchmarks. To better understand this problem, we analyzed prior work on agentic benchmark pitfalls [87, 35, 80, 62, 46] and 17 widely used agentic benchmarks (Table LABEL:tab:all-benchmarks), such as SWE-bench-Verified [14], GAIA [47], $\tau$-bench [86], and WebArena [93]. Combining insights from the literature with our own experience in developing benchmarks, we identified two major challenges to the validity of benchmark results:
@@
Following prior work on analyzing AI and code benchmarks [65, 10], we formulate our insights into an Agentic Benchmark Checklist (ABC) to assist benchmark developers and users in critically designing and assessing agentic benchmarks. Using ABC, we assessed ten popular agentic benchmarks that span the full range of agent capabilities, resulting in seven benchmarks with flaws in outcome validity, seven with issues in task validity, and all with limitations in the result reporting. In addition to the issues found in $\tau$-bench-Airline, some other example issues we found are: (1) an agent can score 100% on SWE-Lancer without resolving any tasks; (2) KernelBench overestimates agents’ capabilities in generating correct kernel functions by 31% in absolute terms due to incomprehensive fuzz testing; (3) WebArena overestimates performance of agents by 5.2% due to various issues in its string matching. To demonstrate ABC’s practical value, we applied it to improve CVE-Bench, a complex, representative cybersecurity benchmark [96]. ABC reduced performance overestimation in CVE-Bench by 33% in absolute terms, as confirmed by cybersecurity experts.
@@ 2 Related Work
Benchmarking of AI Agents. Prior work has proposed agentic benchmarks across various domains, including coding [30, 48, 37, 59], interacting with environments for a predefined target [93, 83, 86], solving math problems [22, 39], and others [11, 89, 47, 74]. These tasks typically emulate real-world challenge resolution, involving non-categorical outputs and multistep execution. Evaluating AI agents in these tasks introduces a more complex design and implementation than traditional benchmarks, including the handling dynamic interactions between an agent and the environment and grading unstructured responses, which increases the difficulty in ensuring rigorous evaluation.
Issues in Evaluating AI Agents. Existing analyses have identified evaluation issues in individual agentic benchmarks [32, 62, 34, 35, 87]. In terms of the outcome validity, Kydlíček and Gandenberger [34] found that implicit assumptions on the answer formats lead to performance underestimation by 5.3%. In addition, Yu et al. [87] found that agents can pass evaluations without generating correct patches for 7.7% of tasks in the SWE-bench-Lite and 5.2% of tasks in the SWE-bench-Verified. In terms of task validity, prior analysis found that the annotation noise in BIRD significantly affects the accuracy of performance evaluation [62, 80]. The rate limit of the websites implemented in WebArena prevented agents from resolving challenges [32]. Furthermore, Lange et al. [35] identified flaws in the grading of KernelBench that allow agents to bypass correctness checks. However, none of them develops an actionable and systematic guideline to assess agentic benchmarks.
@@ 3 Overview
In this section, we present an over of our work. We first introduce a taxonomy of validity issues in agentic benchmarks and the describe the process of our benchmark collection, checklist development, and benchmark assessment. Finally, we release our code [^1] https://github.com/uiuc-kang-lab/agentic-benchmarks and build a website [^2] https://uiuc-kang-lab.github.io/agentic-benchmarks/ for continuous development and future updates.
@@
First, improper task setup can lead to the violation of task validity. For instance, $\tau$-bench includes intentionally unattainable tasks (e.g., making changes to a non-refundable ticket), which agents are supposed to recognize and reject [86]. Yet, a trivial agent that simply returns nothing is considered as a successful completion even though it cannot look up information or interpret ticket rules. Second, failure to rigorously grade unstructured task outcome can break the outcome validity. For example, SWE-bench-Verified judges agent-generated patches by handwritten unit tests [14]. Since such tests can be incomplete or not perfectly sound [94, 87], a patch that passes them may still be wrong. Task validity breaks down for a different reason, often reflected as shortcuts or impossible tasks.
@@
Benchmark Collection. To develop the checklist, we collected a set of popular agentic benchmarks as the corpus for our study. To emphasize common and representative issues, we focused on popular agentic benchmarks used by top AI providers, including OpenAI, Anthropic, Amazon, Meta, Google, xAI, Mistral, and DeepSeek, or those winning awards in peer-reviewed academic conferences. That narrows our focus to a set of 17 agentic benchmarks (Table LABEL:tab:all-benchmarks). We defer the details of our benchmark collection to Appendix B.
Checklist Development. We first reviewed the collected benchmarks and surveyed AI agent evaluation frameworks [1, 50, 44, 45] together with documented issues in agentic benchmarks [32, 62, 34, 35, 87]. We then examined best practices for evaluating unstructured task outcomes in related domains, such as software testing. Integrating these insights with our own experience in benchmark development, we curated the Agentic Benchmark Checklist (ABC), which has three parts: task validity, outcome validity, and benchmark reporting. We provide the source of each checklist items in Appendix C.
Benchmark Assessment. We applied ABC to thoroughly assess ten selected benchmarks (Table 1). We selected these benchmarks from the open-source set in Table LABEL:tab:all-benchmarks, prioritizing their popularity and ensuring all types of agents’ capabilities are covered. We assigned 1 point to each satisfied item and 0 otherwise. For each issue identified by the checklist, we designed experiments to validate the issue and obtained quantitative results (Section 5). We defer detailed assessment results to Appendix D and case studies to Appendix E.
@@ 4.1 Assessing Task Validity
Environment. Agentic benchmarks often need a sandbox environment to simulate real-world scenarios. Implementing and maintaining such environments can be challenging, especially with complex task formulations. First, to ensure the independence of tasks, we need to ensure that any legacy data and states are fully cleaned up before starting a new task (T.4). For example, KernelBench failed to remove ground truth answers from GPU memory, allowing agents to obtain the correct result through out-of-bounds memory access [35]. Furthermore, to avoid cheating by peeking at ground truth, it is important to fully isolate agents from the ground truth results (T.5). Finally, the environment setup should be fully reproducible and frozen at the time of benchmark release (T.6). Relying on dynamic resources, such as continually updated external websites, are not recommended.
@@ 4.2 Assessing Outcome Validity
In this part of the assessment, we propose practical checks for ensuring the outcome validity of an agentic benchmark (Figure 3). We design these checks based on different type of outcomes and different evaluation methods.
Information Acquisition. To evaluate the capability of AI agents to search, retrieve, integrate, and summarize information, agentic benchmarks formulate tasks as information acquisition queries [86, 25, 89, 93]. Depending on task requirements, benchmarks use various schemes for evaluating agent’s textual responses, including whole string matching [89], substring matching [86, 93], and LLM-as-a-judge [93, 25].
@@
    Whole String Matching directly compares $R_{i}^{(A)}$ and $R_{i}^{(G)}$. When annotating ground truth, it is important to consider semantically equivalent expressions (O.a.1) or redundant words (O.a.2). [^3] In practice, users often specify format requirements for AI agents, which narrows the scope of alternative expressions of the ground truth. Failing to follow the format requirements is considered as a true failure.
@@
    Substring Matching evaluates whether $R_{i}^{(A)}$ contains the $R_{i}^{(G)}$. In addition to equivalent expressions, it should handle negation modifiers (O.b.1), such as “not” and “negative.” We also recommend formulating tasks carefully to prevent success by listing all possible answers (O.b.2) or guessing (O.b.3).
@@
    LLM-as-a-Judge uses LLM(s) to emulate human annotators [90, 88, 9, 38, 97]. Previous studies have shown that the accuracy of LLM annotations varies across domains [98]. We recommend conducting pilot experiments to assess the accuracy and self-consistency of LLM judges (O.c.1).
@@ 5.1 Assessment Scores
{tblr}
[…]
@@ Appendix B Details of Benchmark Collection and Selection
[…]
@@
[…]
@@ Appendix C Sources of the Checks in ABC
[…]
@@
O.h.1 Mialon et al. [47] Kydlíček and Gandenberger [34],
@@ Appendix D Assessment Reports
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@ E.2 $\tau$-bench
Qualitative Results. A do-nothing agent that returns immediately can achieve a 38% and 6.0% pass^k or pass@k for any k for Airline and Retail partitions, respectively. A spamming agent that outputs all the data can achieve a 40% and 9.6% pass^k or pass@k for any k for Airline and Retail partitions, respectively.
@@ Appendix F An Example of Rigorous Benchmark Reporting
[…]

### v5

@@ 1 Introduction
Although issues in evaluation rigor can significantly skew evaluation results, they are still frequently overlooked in the current development, deployment, and analysis of agentic benchmarks. To better understand this problem, we analyzed prior work on agentic benchmark pitfalls [87, 35, 80, 62, 46] and 17 widely used agentic benchmarks (Table LABEL:tab:all-benchmarks), such as SWE-bench-Verified [14], GAIA [47], $\tau$-bench [86], and WebArena [93]. Combining insights from the literature with our own experience in developing benchmarks, we identified two major conditions of the validity of benchmark results:
@@
Following prior work on analyzing AI and code benchmarks [65, 10], we formulate our insights into an Agentic Benchmark Checklist (ABC) to assist benchmark developers and users in critically designing and assessing agentic benchmarks. Using ABC, we assessed ten popular agentic benchmarks that span the full range of agent capabilities, resulting in seven benchmarks with flaws in outcome validity, seven with issues in task validity, and all with limitations in the result reporting. In addition to the issues found in $\tau$-bench-Airline, some other example issues we found are: (1) an agent can score 100% on SWE-Lancer [48] without resolving any tasks; (2) KernelBench [59] overestimates agents’ capabilities in generating correct kernel functions by 31% in absolute terms due to incomprehensive fuzz testing; (3) WebArena [93] overestimates performance of agents by 5.2% due to various issues in its string matching. To demonstrate ABC’s practical value, we applied it to improve CVE-Bench, a complex, representative cybersecurity benchmark [96]. ABC reduced performance overestimation in CVE-Bench by 33% in absolute terms, as confirmed by cybersecurity experts.
@@ 2 Related Work
Benchmarking of AI Agents. Prior work has proposed agentic benchmarks across various domains, including coding [30, 48, 37, 59], interacting with environments for a predefined target [93, 83, 86], solving math problems [22, 39], and others [11, 89, 47, 74]. These tasks typically emulate real-world challenge resolution, involving non-categorical outputs and multistep execution. Evaluating AI agents in these tasks introduces a more complex design and implementation than traditional benchmarks, including handling dynamic interactions between an agent and the environment and grading unstructured responses, which increases the difficulty in ensuring rigorous evaluation.
Issues in Evaluating AI Agents. Existing analyses have identified evaluation issues in individual agentic benchmarks [32, 62, 34, 35, 87]. In terms of the outcome validity, Kydl´ıček and Gandenberger [34] found that implicit assumptions on the answer formats lead to performance underestimation by 5.3%. Yu et al. [87] found that agents can pass evaluations without generating correct patches for 7.7% of tasks in the SWE-bench-Lite and 5.2% of tasks in the SWE-bench-Verified. In addition, prior analysis found that the annotation noise in BIRD significantly affects the accuracy of performance evaluation [62, 80]. In terms of task validity, the rate limit of the websites implemented in WebArena prevented agents from resolving challenges [32]. Furthermore, Lange et al. [35] identified flaws in the grading of KernelBench that allow agents to bypass correctness checks. However, none of them develops an actionable and systematic guideline to assess agentic benchmarks.
@@ 3 Overview
In this section, we present an overview of our work. We first introduce a taxonomy of validity issues in agentic benchmarks and then describe the process of our benchmark collection, checklist development, and benchmark assessment. Finally, we release our code [^1] https://github.com/uiuc-kang-lab/agentic-benchmarks and build a website [^2] https://uiuc-kang-lab.github.io/agentic-benchmarks/ for continuous development and future updates.
@@
First, improper task setup can lead to the violation of task validity. For instance, $\tau$-bench includes intentionally unattainable tasks (e.g., making changes to a non-refundable ticket), which agents are supposed to recognize and reject [86]. Yet, a trivial agent that simply returns nothing is considered a successful completion even though it cannot look up information or interpret ticket rules. Second, failure to rigorously grade unstructured task outcome can break outcome validity. For example, SWE-bench-Verified judges agent-generated patches by handwritten unit tests [14]. Since such tests can be incomplete or not perfectly sound [94, 87], a patch that passes them may still be wrong. Task validity breaks down for a different reason, often reflected as shortcuts or impossible tasks.
@@
Benchmark Collection. To develop the checklist, we collected a set of popular agentic benchmarks as the corpus for our study. To emphasize common and representative issues, we focused on popular agentic benchmarks used by top AI providers, including OpenAI, Anthropic, Amazon, Meta, Google, xAI, Mistral, and DeepSeek, or those winning awards in peer-reviewed academic conferences. This narrows our focus to a set of 17 agentic benchmarks (Table LABEL:tab:all-benchmarks). We defer the details of our benchmark collection to Appendix B.
Checklist Development. We first reviewed the collected benchmarks and surveyed AI agent evaluation frameworks [1, 50, 44, 45] together with documented issues in agentic benchmarks [32, 62, 34, 35, 87]. We then examined best practices for evaluating unstructured task outcomes in related domains, such as software testing. Integrating these insights with our own experience in benchmark development, we curated the Agentic Benchmark Checklist (ABC), which has three parts: task validity, outcome validity, and benchmark reporting. We provide the source of each checklist item in Appendix C.
Benchmark Assessment. We applied ABC to thoroughly assess ten selected benchmarks (Table 1). We selected these benchmarks from the open-source set in Table LABEL:tab:all-benchmarks, prioritizing their popularity and ensuring all types of agent capabilities are covered. We assigned 1 point to each satisfied item and 0 otherwise. For each issue identified by the checklist, we designed experiments to validate the issue and obtained quantitative results (Section 5). We defer detailed assessment results to Appendix D and case studies to Appendix E.
@@ 4.1 Assessing Task Validity
Environment. Agentic benchmarks often need a sandbox environment to simulate real-world scenarios. Implementing and maintaining such environments can be challenging, especially with complex task formulations. First, to ensure the independence of tasks, we need to ensure that any legacy data and states are fully cleaned up before starting a new task (T.4). For example, KernelBench failed to remove ground truth answers from GPU memory, allowing agents to obtain the correct result through out-of-bounds memory access [35]. Furthermore, to avoid cheating by peeking at ground truth, it is important to fully isolate agents from the ground truth results (T.5). Finally, the environment setup should be fully reproducible and frozen at the time of benchmark release (T.6). Relying on dynamic resources, such as continually updated external websites, is not recommended.
@@ 4.2 Assessing Outcome Validity
In this part of the assessment, we propose practical checks for ensuring the outcome validity of an agentic benchmark (Figure 3). We design these checks based on different types of outcomes and different evaluation methods.
Information Acquisition. To evaluate the capability of AI agents to search, retrieve, integrate, and summarize information, agentic benchmarks formulate tasks as information acquisition queries [86, 25, 89, 93]. Depending on task requirements, benchmarks use various schemes for evaluating agents’ textual responses, including whole string matching [89], substring matching [86, 93], and LLM-as-a-judge [93, 25].
@@
    Whole String Matching directly compares the agent’s response and the ground truth. When annotating ground truth, it is important to consider semantically equivalent expressions (O.a.1) or redundant words (O.a.2). [^3] In practice, users often specify format requirements for AI agents, which narrows the scope of alternative expressions of the ground truth. Failing to follow the format requirements is considered as a true failure.
@@
    Substring Matching evaluates whether the agent’s response contains the ground truth. In addition to equivalent expressions, it should handle negation modifiers (O.b.1), such as “not” and “negative.” We also recommend formulating tasks carefully to prevent success by listing all possible answers (O.b.2) or guessing (O.b.3).
@@
    LLM-as-a-Judge uses LLMs to emulate human annotators [90, 88, 9, 38, 97]. Previous studies have shown that the accuracy of LLM annotations varies across domains [98]. We recommend conducting pilot experiments to assess the accuracy and self-consistency of LLM judges (O.c.1).
@@ 5.1 Assessment Scores
Benchmark Evaluated Capability Evaluation Design
SWE-bench [30] Software Engineering Unit Testing
@@ Appendix B Details of Benchmark Collection and Selection
[ caption = Benchmarks used by major AI providers between 1 January 2024 and 18 March 2025. Duplicate benchmarks are listed only once., label = tab:all-bench ]colspec=Q[c]Q[l]Q[l]Q[l],row1 = font=, rowodd[2] = bg=gray!25Benchmark Used by Source Agentic
SimpleQA OpenAI Introducing GPT-4.5 [57]
@@
[ caption = Collected agentic benchmarks. Assessed benchmarks are highlighted in blue., label = tab:all-benchmarks ]colspec=Q[c]Q[c]Q[l],row1 = font=, rowodd[2] = bg=gray!25Benchmark Evaluated Capability Evaluation Design
SWE-bench [30] Software Engineering Unit Testing
@@ Appendix C Sources of the Checks in ABC
[ caption = Sources of items in ABC, label = tab:sources ]colspec=Q[c]Q[l]Q[l]Q[l],row1 = font=, rowodd[2] = bg=gray!25Question Existing Best Practice Lessons Learned Domain-Specific Suggestions
O.a.1 Mialon et al. [47],
@@
O.h.1 Mialon et al. [47] Kydl´ıček and Gandenberger [34],
@@ Appendix D Assessment Reports
[ caption = Assessment Report of SWE-Bench-Lancer (paper, code), label = tab:sources ]colspec=Q[c]Q[l]Q[l]Q[l],row1 = font=, rowodd[2] = bg=gray!25Check Score Reason
O.d.1 1 As discussed in Section 1 of the paper, the benchmark uses a set of test cases that are verified for correctness
@@
{longtblr}[ caption = Assessment Report of Bird-Bench (paper, code), label = tab:sources ]colspec=Q[c]Q[l]Q[l]Q[l],row1 = font=, rowodd[2] = bg=gray!25Check Score Reason
O.d.1 1 As discussed in Section 3.4 of the paper, the validity of the database is verified by executing the ground-truth
@@
{longtblr}[ caption = Assessment Report of CyBench (paper, code), label = tab:sources ]colspec=Q[c]Q[l]Q[l]Q[l],row1 = font=, rowodd[2] = bg=gray!25Check Score Reason
O.h.1 1 The specific format required for the answer is provided in the task description.
@@
{longtblr}[ caption = Assessment Report of SWE-Bench-Verified (paper, code), label = tab:sources ]colspec=Q[c]Q[l]Q[l]Q[l],row1 = font=, rowodd[2] = bg=gray!25Check Score Reason
O.d.1 1 Test cases are directly taken from GitHub repositories, and the paper does not mention any verification process.
@@
{longtblr}[ caption = Assessment Report of $\ tau$-Bench (paper, code), label = tab:sources ]colspec=Q[c]Q[l]Q[l]Q[l],row1 = font=, rowodd[2] = bg=gray!25Check Score Reason
O.a.1 1 The benchmark uses minimal expressions for substring matching, which is robust to variations in the input.
@@
{longtblr}[ caption = Assessment Report of MLE-Bench (paper, code), label = tab:sources ]colspec=Q[c]Q[l]Q[l]Q[l],row1 = font=, rowodd[2] = bg=gray!25Check Score Reason
O.I.1 1 As described in Section 2.2, the benchmark uses leaderboard positions as a metric, which is not easily exploitable.
@@
{longtblr}[ caption = Assessment Report of WebArena (paper, code), label = tab:sources ]colspec=Q[c]Q[l]Q[l]Q[l],row1 = font=, rowodd[2] = bg=gray!25Check Score Reason
O.a.1 1 As discussed in Section 3.2 of the paper, the benchmark expects the response to follow a standardized format,
@@
{longtblr}[ caption = Assessment Report of GAIA (paper, code), label = tab:sources ]colspec=Q[c]Q[l]Q[l]Q[l],row1 = font=, rowodd[2] = bg=gray!25Check Score Reason
O.h.1 1 As discussed in Section 3.2 of the paper, the specific format required for the answer is provided in the task
@@
{longtblr}[ caption = Assessment Report of OSWorld (paper, code), label = tab:sources ]colspec=Q[c]Q[l]Q[l]Q[l],row1 = font=, rowodd[2] = bg=gray!25Check Score Reason
O.g.1 1 As discussed in Section 3.2 of the paper, the ground truth is verified to include all states that can be achieved
@@
{longtblr}[ caption = Assessment Report of KernelBench (paper, code), label = tab:sources ]colspec=Q[c]Q[l]Q[l]Q[l],row1 = font=, rowodd[2] = bg=gray!25Check Score Reason
O.e.1 0 The fuzzer does not address potential edge cases, such as empty inputs.
@@ E.2 $\tau$-bench
Qualitative Results. A do-nothing agent that returns immediately can achieve a 38% and 6.0% passˆk or pass@k for any k for Airline and Retail partitions, respectively. A spamming agent that outputs all the data can achieve a 40% and 9.6% passˆk or pass@k for any k for Airline and Retail partitions, respectively.
@@ Appendix F An Example of Rigorous Benchmark Reporting
[ caption = Modified Leaderboards of BIRD [37] with Confidence Intervals., label = tab:bird-data ]colspec=Q[c]Q[r]Q[r]Q[r]Q[r],row1 = font=, rowodd[2] = bg=gray!25Method Dev. Accuracy (%) Confidence Interval Original Rank Possible Rank
CHASE-SQL + Gemini74.9[66.8, 71.4]11-13